OpenStudy (anonymous):

v + 3/5 = 2 1/3 please help

OpenStudy (karatechopper):

first put 2 1/3 in an improper fraction for me..

OpenStudy (karatechopper):

improper fraction would be 7/3

OpenStudy (karatechopper):

next you need to get a common denominator..so what r some common denominators?

OpenStudy (anonymous):

\[v + \frac{3}{5} = 2 \frac{1}{3}\]\[v + \frac{3}{5} = \frac{7}{3}\]\[v = \frac{7}{3}- \frac{3}{5}\]\[v = \frac{7*5}{15}- \frac{3*3}{15}\]\[v = \frac{35}{15}- \frac{9}{15}\]\[v = \frac{35-9}{15}\]\[v = \frac{26}{15}\]